OpenStudy (anonymous):

You are painting the walls in a room of a new house. your task is to determine the cost of paint for the room with given dimensions.

OpenStudy (anonymous):

Step 1: Fin the area of the walls you need to paint

OpenStudy (anonymous):

@Peaches15 @peachpi @izuru @phi @dan815

OpenStudy (anonymous):

Wall A's area is easy to find. Just multiply the length and the width. For the others, try and divide the rectangles into small pieces. With those small pieces, find the area as well. Then, add the areas to find the total area that needs to be covered in paint.

OpenStudy (anonymous):

@izuru Wall 1 would be 162ft right? And wall 3 would be 134ft

OpenStudy (anonymous):

But I don't know how to get wall 2 and 4

OpenStudy (anonymous):

@rhr12

OpenStudy (anonymous):

@DanJS

OpenStudy (rhr12):

I don't get the question. sorry.

OpenStudy (anonymous):

I need to find the area of the walls in the picture

OpenStudy (danjs):

wall 1 A = 9 * 18

OpenStudy (anonymous):

Its 162ft right? And wall 3 is 134ft

OpenStudy (danjs):

Wall 2 A = 21*9 - 7*5

OpenStudy (anonymous):

So 154ft

OpenStudy (danjs):

wall 3 A = 18*9 - 7*4

OpenStudy (danjs):

Wall 4 A = 21*9 - 4*3

OpenStudy (anonymous):

Thank you @DanJS can you help me with another question?

OpenStudy (danjs):

ok

OpenStudy (anonymous):

That was step 1 the next step says what is the total area that we need to cover with paint? Do I just add all of those areas together then? Or is there something else I feel like im over analyzing this. @DanJS